OpenStudy (anonymous):

verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x)(tan x cos x- cot x cos x)= 1- 2 cos^2x

OpenStudy (anonymous):

\[L.H.S=\sin x \left( \tan x \cos x-\cot x \cos x \right)\] \[=\sin x \left( \frac{ \sin x }{ \cos x }*\cos x-\frac{ \cos x }{ \sin x }*\cos x\right)=\sin ^2x-\cos ^2x=1-\cos ^2x-\cos ^2x=?\]

OpenStudy (anonymous):

-cos^2x-cos^2x= -2cos^2x so it would equal 1-2cos^2x..Thank you!!! @surjithayer

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you help me with some more? I have three more to go..

OpenStudy (anonymous):

@surjithayer